# Complements of discriminants of real parabolic function singularities. II

### v4

@@ 5. Petrovskii lacunas
The above enumeration of the connected components of the discriminant complements at parabolic singularities provides a complete list of local Petrovskii lacunas at the corresponding singular points of wavefronts of hyperbolic PDEs. In this theory, generating families of functions play a crucial role. Wave functions are defined on the parameter spaces of these families via integrals of special differential forms along certain cycles in the zero-level sets of the corresponding functions. Wavefronts are the discriminant sets of these families. Outside of the wavefront, the integration cycles and integrand forms depend regularly on the parameters, so that the wave function is regular as well. However, over the discriminant sets of parameters, the level sets degenerate and the wave function may become irregular. The type of irregularity of a wave function near a wavefront point is characterized by the asymptotics of the wave function as the non-discriminant parameter values approach the point. The type of this asymptotic behavior primarily depends on the singularity class of the corresponding discriminant function and on the homology class of the integration cycle. These homology classes depend on the connected components of the wavefront complement from which we approach the wavefront point. Such a component is called a local lacuna if the wave function can be continued from it to a regular function on entire neighborhood of the wavefront point.
@@
In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity for even $N$ and $i_{+}$ is component $(a)$ of Fig. 1. The two lacunas of the $X_{9}^{+}$ singularity for odd $N$ and even $i_{+}$ are the components $(a)$ and $(b)$ of the same figure. The lacunas for the $X_{9}^{-}$ singularities are obtained from the $X_{9}^{+}$ singularities by multiplying of all functions by 1-1. The four lacunas at the $X_{9}^{2}$ singularity with even $N$ and odd $i_{+}$ are shown in picture $(i)$ of Fig. 3, or in its rotation by the angle $\frac{\pi}{2}$, or in the pictures obtained by applying the operation (17) to these two. The unique lacuna at the $J_{10}^{3}$ singularity with even $N$ and odd $i_{+}$ is shown in picture $(m)$ of Fig. 5. The two local lacunas of the $P_{8}^{1}$ singularity with odd $N$ and even $i_{+}$ are the components with $\mbox{Ind}=1$. Three local lacunas of the $P_{8}^{2}$ singularity with odd $N$ and even $i_{+}$ are also the components with $\mbox{Ind}=1$. The only new local lacuna found is the component with $\mbox{Card}=262$ for $P_{8}^{2}$ singularity.

### v5

@@ 5. Petrovskii lacunas
The above enumeration of the connected components of the discriminant complements at parabolic singularities provides a complete list of local Petrovskii lacunas at the corresponding singular points of wavefronts of hyperbolic PDEs. In this theory, generating families of functions play a crucial role. Wave functions are defined on the parameter spaces of these families via integrals of special differential forms along certain cycles in the zero-level sets of the corresponding functions. Wavefronts are the discriminant sets of these families. Outside of the wavefront, the integration cycles and integrand forms depend regularly on the parameters, so that the wave function is regular as well. However, over the discriminant sets of parameters, the zero-level sets degenerate and the wave function may become irregular. The type of irregularity of a wave function near a wavefront point is characterized by the asymptotics of the wave function as the non-discriminant parameter values approach the point. The type of this asymptotic behavior primarily depends on the singularity class of the corresponding discriminant function and on the homology class of the integration cycle. These homology classes depend on the connected components of the wavefront complement from which we approach the wavefront point. Such a component is called a local lacuna if the wave function can be continued from it to a regular function on entire neighborhood of the wavefront point.
@@
In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity for even $N$ and $i_{+}$ is component $(a)$ of Fig. 1. The two lacunas of the $X_{9}^{+}$ singularity for odd $N$ and even $i_{+}$ are the components $(a)$ and $(b)$ of the same figure. The lacunas for the $X_{9}^{-}$ singularities are obtained from these of $X_{9}^{+}$ singularities by multiplying of all functions by 1-1. The four lacunas at the $X_{9}^{2}$ singularity with even $N$ and odd $i_{+}$ are shown in picture $(i)$ of Fig. 3, or in its rotation by the angle $\frac{\pi}{2}$, or in the pictures obtained by applying the operation (17) to these two. The unique lacuna at the $J_{10}^{3}$ singularity with even $N$ and odd $i_{+}$ is shown in picture $(m)$ of Fig. 5. The two local lacunas of the $P_{8}^{1}$ singularity with odd $N$ and even $i_{+}$ are the components with $\mbox{Ind}=1$. Three local lacunas of the $P_{8}^{2}$ singularity with odd $N$ and even $i_{+}$ are also the components with $\mbox{Ind}=1$. The only new local lacuna found is the component with $\mbox{Card}=262$ for $P_{8}^{2}$ singularity.